OpenStudy (anonymous):

Vectors: question below only 7b. I need explanation why the answer is k is greater than or equal to zero

OpenStudy (anonymous):

answer: \(k \ge 0\)

OpenStudy (amistre64):

\[|a||b|~cos\alpha=a\cdot b\] |a| = sqrt(2), |b| = k, a.b = k

OpenStudy (amistre64):

but the real question is in the fact that (0,0) is 90 degrees to all other vectors, and cannot possibly be 45 degree with 1,1

OpenStudy (anonymous):

why it can't be 45 degrees with 1,1?

OpenStudy (amistre64):



OpenStudy (amistre64):

0,0 dot 1,1 = 0, and when a dot product is zero, the vectors are 90 degrees

OpenStudy (amistre64):

consider this in 3 dimensions: 1,1,0 0,k,0 and 0,0,1 when k=0, all we are left with is 1,1,0 and 0,0,1

OpenStudy (anonymous):

sorry, i 'm still trying to understand this :/ where did you get 0,0,1?

OpenStudy (amistre64):

its a vector that is perpendicular to all other vectors in the xy plane; the zero vector in they xy plane is actually a part of the 0,0,1 vector, or line, in xyz space

OpenStudy (anonymous):

oh ok :)

OpenStudy (phi):

k=0 seems problematic. But here is what they are thinking: \[ a \cdot b = |a| |b| \cos \theta \] \[ |a| = \sqrt{2} , |b|= |k| , \cos \theta= \frac{1}{\sqrt{2}}\] \[ a \cdot b = k = \sqrt{2} \ |k| \frac{1}{\sqrt{2}} \\ k= |k|\] k≥0 is the solution

OpenStudy (anonymous):

to clrify the picture with @amistre64 permition:  the angle will always be 45º if k is positive.